OpenStudy (anonymous):

prove the identity (show steps): (tan(x)+1)^2=1+2sin(x)cos(x)/cos^2(x)

OpenStudy (anonymous):

Left hand side =[tan(x)+1]^2 =tan^2(x)+2tan(x)+1 =sin^2(x)/cos^2(x)+2sin(x)/cos(x)+1 =[sin^2(x)+2sin(x)cos(x)]/cos^2(x)+1 not=Right hand side. so please check your question

OpenStudy (anonymous):

wolfram alpha says its true, but wont explain why.

OpenStudy (anonymous):

either you entered it incorrectly into wolfram, or you might want to recheck the equation.... as written, this is not an identity: (tan(x)+1)^2=1+2sin(x)cos(x)/cos^2(x) let x=pi/4.... left side = 4 right side = 3

OpenStudy (anonymous):

i checked it, it is exactly as written. still says true.

OpenStudy (anonymous):

i copy/paste from your problem into wolfram... where does it say it's true? http://www.wolframalpha.com/input/?i=%28tan%28x%29%2B1%29%5E2%3D1%2B2sin%28x%29cos%28x%29%2Fcos%5E2%28x%29

OpenStudy (anonymous):

you missed some parentheses. (tan(x)+1)^2=(1+2sin(x)cos(x))/cos^2(x) http://www.wolframalpha.com/input/?i=%28tan%28x%29%2B1%29^2%3D%281%2B2sin%28x%29cos%28x%29%29%2Fcos^2%28x%29

OpenStudy (anonymous):

correction, I missed some, sorry.

OpenStudy (anonymous):

hmmm... that's prolly why the first helper said this is not an identity also.... that's not what's in your original post....

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

ok.... is see now.... so that last one IS the identity in quesion, right? (tan(x)+1)^2=(1+2sin(x)cos(x))/cos^2(x) ???

OpenStudy (anonymous):

yes, except you cut off the last )

OpenStudy (anonymous):

or not. this is confusing.

OpenStudy (anonymous):

that is the right one

OpenStudy (anonymous):

just to be clear, it's this: \(\large (tanx+1)^2=\frac{1+2sinxcosx}{cos^2x} \)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

looks like putting everything in terms of sine/cosine would be best here... since the right side is basically in terms of sine/cosine, let's see if we can simplify that side first.... working on the right side, can you simplify it any further?

OpenStudy (anonymous):

would making it (1+sin(2x)/cos^2(x) help?

OpenStudy (anonymous):

i wouldn't go that direction.... when proving identities, i try to get everything in terms of sinx, cosx, etc. instead of sin(2x), cos(2x), etc... this is what i'd do: \(\large \frac{1+2sinxcosx}{cos^2x}=\frac{1}{cos^2x}+\frac{2sinxcosx}{cos^2x}= \) ????

OpenStudy (anonymous):

can u simplify that last step?

OpenStudy (anonymous):

yes, but how did you get that from the original?

OpenStudy (anonymous):

plain algebra: \(\large \frac{a+b}{c}=\frac{a}{c}+\frac{b}{c} \)

OpenStudy (anonymous):

i mean from (tan(x)+1)^2?

OpenStudy (anonymous):

we'll get to the left side soon.... we're working ONLY with the right side for now....

OpenStudy (anonymous):

2tan(x)

OpenStudy (anonymous):

1/cos^2(x)+2tan(x)

OpenStudy (anonymous):

do you see how i got to this point: \(\large \frac{1+2sinx{cosx}}{{cos^2x}}=\frac{1}{cos^2x}+\frac{2sinx\cancel{cosx}}{\cancel{cos^2x}^{cosx}}=sec^2x+2tanx \)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok.... good.... that's as far as we can get on the right side... agreed?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now the left side is easy.....

OpenStudy (anonymous):

expand \(\large (tanx+1)^2 \) = ?????

OpenStudy (anonymous):

i see now, thanks

OpenStudy (anonymous):

sure??? ok....

OpenStudy (anonymous):

tan^2(x)+1+2tan(x) tan^2(x)+1=sec^2(x) 2tan(x)+sec^2(x)

OpenStudy (anonymous):

nicely done!!! :)

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

yw... :)